OpenStudy (dmezzullo):

Evaluate the following expression using the values given. Find 3x - y - 3z if x = -2, y = 1, and z = -2. -13 -1 1 13

OpenStudy (anonymous):

Just substitute the numbers: 3(-2) - 1 - 3(-2) Can you solve from here?

OpenStudy (dmezzullo):

no

OpenStudy (dmezzullo):

yes

OpenStudy (anonymous):

yes u can its simple math lol

OpenStudy (anonymous):

Tell what the answer would be

OpenStudy (anonymous):

3 x (-2) = -6 - 1 - 6 So -6 -1 = -7 and -7 - 6 = - 13 :)

OpenStudy (dmezzullo):

-13?